# Causal Network Recovery in Perturb-seq Experiments Using Proxy and Instrumental Variables

### v1

@@ 6.2 Modular structure of inter-chromosomal causal gene networks
Figure 6: In- and out-degree investigation of the inter-chromosomal causal gene network. a,b. Histograms of gene in-degree and out-degree in the DAG. c. Scatterplot summarizing gene in-degree (y-axis) versus out-degree (x-axis), where genes labeled in red (blue) have larger in-degree (out-degree). d. Outer gray pie chart summarizes how many TFs (listed in ENCODE encode2012integrated) in the initial set of analyzed genes were within the DAG with at least one edge. The inner pie chart summarizes the number of TFs with greater out-degree than in-degree (blue) and vice versa (red) among the 18 TFs present in the DAG. e,f. Top five Gene Ontology terms for in-hubs (e) and out-hubs (f).
@@
We examined the structure of the inferred inter-chromosomal network by applying Louvain clustering with edge z-scores (14) as weights, yielding 10 modules annotated by their top three Gene Ontology (GO) terms (Fig. 5a). The clusters form distinct and also coherent functional modules dominated by core cellular processes, consistent with essential-gene categories in replogle2022mapping. Notably: (i) cytoplasmic translation, represented by two (green) modules enriched for ribosomal proteins (RPL/RPS), reflecting the translational demands of proliferative K562 cells (slavov2015differential, khajuria2018ribosome); (ii) mitochondrial translation and electron transport, represented by a (red) module enriched for mitochondrial ribosomal proteins (MRPL/MRPS), which couple mitochondrial translation to respiratory-complex assembly (greber2016structure); and (iii) negative regulation of organelle organization and regulation of chromosome segregation, represented by a (light purple) module enriched for BUB and SMC regulators, consistent with the role of BUB family checkpoint proteins in ensuring proper kinetochore attachment (musacchio2017molecular) and with the spindle and segregation defects observed in K562 cells (martinez2016subpopulation).
@@
The topological structures learned from ARGEN also provide compelling results. First, the hierarchical structure of the DAG is skewed to the right (Fig. 5b), indicating that the DAG reveals a shallow hierarchical architecture. This suggests that most genes act as either immediate or near-direct regulators with regulatory signal rarely propagating through deep chains, supported by previous findings that essential processes form compact regulatory modules with more direct regulators than the ones with cascading effects (alon2007network). In addition, a summary of the DAG with Katz centrality indicates that there are a few hub nodes with high centrality (Fig. 5c). We specifically examined the in and out degree distributions separately (Fig. 6a,b). As in centrality summary, the degree distributions were highly skewed to the right.
@@ 6.3 In- and out-degree distributions of the causal networks
Furthermore, we observed that majorities (310 genes) of the nodes were with greater out degree, while a few outlying genes were with large in-degree (Fig. 6c), showing alignment with the previously reported regulatory network structure obtained from a large scale ENCODE datasets across five different cell lines involving K562 (gerstein2012architecture). Interestingly, among these 400 genes, there were initially 19 transcription factors, essential regulators of genes, and 17 of them were with greater out-degree than in-degree (Fig. 6d), further providing biological validity of the ARGEN DAGs.
Several genes, including MIPEP, HARS2, IPO13, MBTPS2, and ABHD12, emerged as high in-degree sink nodes in the inferred DAG, a pattern that is biologically consistent with their roles as downstream integrators rather than upstream regulators. MIPEP and HARS2 genes perform critical roles in mitochondrial maintenance, a highly sensitive process that must adapt to changing cellular energy demands and stress conditions (Chew1997MIPEP, Diodato2014mtARS). IPO13, a bi-directional nuclear transport receptor, similarly integrates signals arising from changes in transcriptional demand and RNA-processing stress (Mingot2001). MBTPS2, the Site-2 protease required for ATF6 and SREBP activation, is transcriptionally induced under ER stress and lipid-homeostasis disruption (ye2000ATF6cleavage). ABHD12 is a lipid-metabolizing enzyme with an important role in immune regulation and brain function (Kamat2015ABHD). The elevated in-degree reflects downstream responsiveness to diverse cellular perturbations rather than direct regulatory influence, providing an internal validation of the inferred causal network structure. A high degree of regulation allows the cell to maintain the correct levels of these essential proteins under varying conditions, ensuring the robustness and stability of vital cellular processes.
Finally, as an additional exploration, we focused on 20 In-hub and 20 Out-hub genes, defined as those with the largest in-degree and out-degree, respectively (Fig. 5c). Separate GO analyses for each group revealed distinct functionalities: In-hubs were enriched for cytoplasmic translation and cellular amide metabolic process, whereas Out-hubs were enriched for ribosome biogenesis and RNA processing. This division reflects established biology because translation-related genes are highly dosage-sensitive and therefore tightly regulated (torres2007effects), while ribosome biogenesis and RNA-processing genes function as upstream controllers of growth and transcriptional output (lempiainen2009growth).
@@ 7 Discussion
Several extensions merit consideration. First, our causal model assumes that perturbation effects propagate through steady-state transcriptional responses, whereas some regulatory interactions may operate at rapid, post-transcriptional, or translational timescales. Incorporating time-resolved single-cell perturbation data, once more broadly available, may improve resolution of dynamic regulatory mechanisms. Second, our framework is designed to provide an accurate causal gene network over the set of genes that are intervened in Perturb-seq experiments. While this is practically meaningful given that perturbed genes are typically chosen for biological relevance, extending the framework to partial intervention settings and characterizing the resulting identifiable DAGs remain important directions for future work. Third, because the same control cells are reused across genes in our two-step procedure, the resulting edge-level p-values may exhibit mild dependence. Developing online FDR control methods robust to correlated p-values could further strengthen inference in this setting. Finally, although we analyzed a well-characterized essential-gene CRISPRi screen, extending the method to more heterogeneous or context-specific regulatory settings such as developmental systems or tumor microenvironments will require further methodological refinements and experimental validation.

### v2

@@ 6.2 Modular structure of inter-chromosomal causal gene networks
Figure 6: In- and out-degree investigation of the inter-chromosomal causal gene network. a,b. Histograms of gene in-degree and out-degree in the DAG. c. Scatterplot summarizing gene in-degree (y-axis) versus out-degree (x-axis), where genes labeled in red (blue) have larger in-degree (out-degree). d. Outer gray pie chart summarizes how many TFs (listed in ENCODE Consortium and others (2012)) in the initial set of analyzed genes were within the DAG with at least one edge. The inner pie chart summarizes the number of TFs with greater out-degree than in-degree (blue) and vice versa (red) among the 18 TFs present in the DAG. e,f. Top five Gene Ontology terms for in-hubs (e) and out-hubs (f).
@@
We examined the structure of the inferred inter-chromosomal network by applying Louvain clustering with edge z-scores (13) as weights, yielding 10 modules annotated by their top three Gene Ontology (GO) terms (Fig. 5a). The clusters form distinct and also coherent functional modules dominated by core cellular processes, consistent with essential-gene categories in Replogle et al. (2022). Notably: (i) cytoplasmic translation, represented by two (green) modules enriched for ribosomal proteins (RPL/RPS), reflecting the translational demands of proliferative K562 cells (Slavov et al. 2015, Khajuria et al. 2018); (ii) mitochondrial translation and electron transport, represented by a (red) module enriched for mitochondrial ribosomal proteins (MRPL/MRPS), which couple mitochondrial translation to respiratory-complex assembly (Greber and Ban 2016); and (iii) negative regulation of organelle organization and regulation of chromosome segregation, represented by a (light purple) module enriched for BUB and SMC regulators, consistent with the role of BUB family checkpoint proteins in ensuring proper kinetochore attachment (Musacchio and Desai 2017) and with the spindle and segregation defects observed in K562 cells (Martinez-Castillo et al. 2016).
@@
The topological structures learned from ARGEN also provide compelling results. First, the hierarchical structure of the DAG is skewed to the right (Fig. 5b), indicating that the DAG reveals a shallow hierarchical architecture. This suggests that most genes act as either immediate or near-direct regulators with regulatory signal rarely propagating through deep chains, supported by previous findings that essential processes form compact regulatory modules with more direct regulators than the ones with cascading effects (Alon 2007). In addition, a summary of the DAG with Katz centrality indicates that there are a few hub nodes with high centrality (Fig. 5c). Examining the in- and out-degree distributions separately (Fig. 6a,b) yields a similar right-skewed pattern.
@@ 6.3 In- and out-degree distributions of the causal networks
Furthermore, we observed that majorities (310 genes) of the nodes were with greater out-degree, while a few outlying genes were with large in-degree (Fig. 6c), showing alignment with the previously reported regulatory network structure obtained from a large scale ENCODE datasets across five different cell lines involving K562 (Gerstein et al. 2012). Interestingly, among these 400 genes, there were initially 19 transcription factors, essential regulators of genes, and 17 of them were with greater out-degree than in-degree (Fig. 6d), further providing biological validity of the ARGEN DAGs.
Several genes, including MIPEP, HARS2, IPO13, MBTPS2, and ABHD12, emerged as high in-degree sink nodes in the inferred DAG, a pattern that is biologically consistent with their roles as downstream integrators rather than upstream regulators. MIPEP and HARS2 genes perform critical roles in mitochondrial maintenance, a highly sensitive process that must adapt to changing cellular energy demands and stress conditions (Chew and others 1997, Diodato et al. 2014). IPO13, a bi-directional nuclear transport receptor, similarly integrates signals arising from changes in transcriptional demand and RNA-processing stress (Mingot et al. 2001). MBTPS2, the Site-2 protease required for ATF6 and SREBP activation, is transcriptionally induced under ER stress and lipid-homeostasis disruption (Ye et al. 2000). ABHD12 is a lipid-metabolizing enzyme with an important role in immune regulation (Kamat et al. 2015). The elevated in-degree reflects downstream responsiveness to diverse cellular perturbations rather than direct regulatory influence, providing an internal validation of the inferred causal network structure. A high degree of regulation allows the cell to maintain the correct levels of these essential proteins under varying conditions, ensuring the robustness and stability of vital cellular processes.
Finally, as an additional exploration, we focused on 20 in-hub and 20 out-hub genes, defined as those with the largest in-degree and out-degree, respectively (Fig. 5c). Separate GO analyses for each group revealed distinct functionalities: In-hubs were enriched for cytoplasmic translation and cellular amide metabolic process, whereas Out-hubs were enriched for ribosome biogenesis and RNA processing. This division reflects established biology because translation-related genes are highly dosage-sensitive and therefore tightly regulated (Torres et al. 2007), while ribosome biogenesis and RNA-processing genes function as upstream controllers of growth and transcriptional output (Lempiäinen and Shore 2009).
@@ 7 Discussion
Several extensions merit consideration. First, our causal model assumes that perturbation effects propagate through steady-state transcriptional responses, whereas some regulatory interactions may operate at rapid, post-transcriptional, or translational timescales. Incorporating time-resolved single-cell perturbation data, once more broadly available, may improve resolution of dynamic regulatory mechanisms. Second, our framework is designed to provide an accurate causal network over the set of genes that are intervened in Perturb-seq experiments. While this is practically meaningful given that perturbed genes are typically chosen for biological relevance, extending the framework to partial intervention settings and characterizing the resulting identifiable DAGs remain important directions. Third, because the same control cells are reused across genes in our procedure, the resulting edge-level p-values may exhibit mild dependence. Developing online FDR control methods robust to correlated p-values could further strengthen inference. Finally, although we analyzed a well-characterized essential-gene CRISPRi screen, extending the method to more context-specific regulatory settings such as developmental systems or tumor microenvironments will require further methodological refinements and experimental validation.
@@ Acknowledgements
Data Availability Statement
Data and R codes to reproduce the results are available in the ARGEN Github repository.
Supplementary Materials
The online Supplementary Materials include details on identifiability of the causal network, additional Lemmas and Theorems and all the proofs, main simulation designs and results, details on data processing and implementation, and additional figures.